OpenStudy (anonymous):

The larger of two consecutive integers is 7 greater than twice the smaller. Find the integers.

OpenStudy (anonymous):

if \(n\) is the smaller, then the larger is \(n+1\) and you know \(n+1= 2n+7\)

OpenStudy (anonymous):

solve \[n+1=2n+7\] \[1=n+7\] \[-6=n\] so the smaller is \(-6\) and the larger is \(-5\)